OpenStudy (anonymous):

Write negative three point two as a mixed number in simplest form. negative three and two-fifths negative three and one-fifth negative three and one-third negative three and one-twentieth

OpenStudy (anonymous):

\[-3.2=-\left(3+\frac{2}{10}\right)=-\left(3+\frac{1}{5}\right) \]" negative three and one-third"